Here's my problem:
If you are 'h' feet above the earth's surface, the distance, 'd' that you can see is approximately 'd' (in miles) = the square root of 3h/2.
Suppose you are on the roof of a building 900ft. tall. How far can you see? (Calculate to 2 decimal places.)
Thanks for your help!
Joanne

Answer by checkley77(12844)

d=sqrt(3h/2)
d=sqrt(3*900/2)
d=sqrt(2700/2)
d=sqrt1350
d=36.74 miles.